The area of a right triangle is 120 sq. in. The longer leg is 4 more than twice the shorter leg. Find the length of each leg. Show all work to receive credit.

Answer by checkley77(12844)

Area=bh/2
b=2h+4
120=(2h+4)h/2
120=h^2+2h/2
120=h^2+2h
h^2+2h-120=0
(h-10)(h+12)
h-10=0
h=10 ans.
b=2*10+4
h=20+4
h=24 ans.
Proof:
120=10*24/2
120=240/2
120=120